OpenStudy (anonymous):

A 59 meter lighthouse spots a distress signal from a sailboat 151 meters from the base of the lighthouse. What is the angle of depression, to the nearest degree, from the lighthouse to the sailboat?

OpenStudy (anonymous):

use trig. height of lighthouse will be ur base or perpendicular...

OpenStudy (anonymous):

Can you extrapolate a bit more?

OpenStudy (dumbcow):

It makes a right triangle, with the line from the top of lighthouse to the boat as the hypotenuse. Find the angle between that line and the lighthouse. Use tan function, tan = opp/adj

OpenStudy (anonymous):

The angle of depression from the lighthouse is the same as the angle of elevation from the sailboat:\[\text{ArcTan}\left[\frac{59}{151}\right]=21^{\circ} \]

OpenStudy (anonymous):

Thank you